please solve this equation
solve t:

5t/12=4-3t

Answer by Maths68(1474)

5t/12=4-3t
Multiply by 12 both sides of the equation
12*(5t/12)=12*(4-3t)
5t=48-36t
5t+36=48
41t=48
Divide by 41 both sides of the equation
41t/41=48/41
t=48/41
Check
Put the value of t in original equation
5t/12=4-3t
5 (48/41)/12=4-3(48/41)
5 (48/41)* 1/12=((4*41)-(3*48)/41
5(48/(41*12))=(164-144)/41
5(4/41)=20/41
20/41=20/41